The nth term of a linear sequence is 12-4n. find
1.first term
2.common difference

Answer by stanbon(75887)

The nth term of a linear sequence is 12-4n. find
1.first term
f(1) = 12-4*1 = 8
-------------
2.common difference
f(2) = 12-4*2) = 4
----
common difference is 4-8 = -4
------------
Cheers,
Stan H.
-----------